Circular No. 6129

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SATURN F. Colas, Bureau des Longitudes, Paris; E. Frappa, Planetarium de

Saint-Etienne; J. Gomez, G.E.A., Barcelona; P. Laques, Observatoire du Pic du Midi; J. Lecacheux, Observatoire de Meudon, DESPA; and M. Tagger, Service d’Astrophysique, Centre d’Etudes de Saclay, report on recent observations at the Pic du Midi 1-m telescope: “During continuous CCD imaging of Saturn on 1994 Nov. 22.8 and Dec. 16.8–17.8 UT under very stable seeing conditions (0′′.6), we convincingly detected several ‘spokes’ on the B ring [see also Sheehan and O’Meara 1993, Sky Tel. 85(1), 20], the rings being then inclined by 8^{o} to our line-of-sight. Our best time resolution was 4 images/min. The reflectance contrast for the features (spokes) in the rings was at maximum in the I band (855 nm), reaching 3 percent on our best images, but they were also detected in R; from our 3.8-hr sequence on Dec. 17 (169 images), their behavior was reminiscent of that detected by both Voyager spacecraft, appearing on that part of the rings exiting the saturnian shadow. The spokes are observed close to the corotation radius (where the keplerian frequency equals the rotation frequency of Saturn and its magnetosphere), with some extension inwards and outwards by ± a few thousands of km. The spokes vanished upon arrival in front of the planet. No significant detection was obtained on the opposite half-ring. The maximum activity was observed when dawn occurred on the rings over 95^{o}–125^{o} in longitude (system III), consistent with the strong peak at 115^{o} observed from Voyager data (cf. Esposito et al. and Mendis et al. 1984, in Saturn, Univ. of Arizona Press). Saturn is currently nearing solar conjunction, and the current ring tilt is 5^{o}.7; on April 12, at a similar elongation after conjunction, the tilt will be only 1^{o}.6 and the spokes difficult or impossible to image through the ring-plane passages.”

RX J0558+53 K. Kasturirangan, Indian Space Research Organisation (ISRO), Ban-

galore, reports: “B. N. Ashoka, T. M. K. Marar, and S. Seetha (ISRO Satellite Centre) conducted high-speed photometric observations on Jan. 8 with the 1-m telescope at Vanna Bappu Observatory, Kavalur, on the recently discovered intermediate polar RX J0558+53. Optical pulsations in white light were unambiguously detected at the reported x-ray period of 272 s. Typical intensity modulation in the light curve was 15 percent. Tentative evidence of the beat period between 272 s and the orbital period of 4.15 hr is also seen in the Fourier transform of the data.”

1995 January 27 Daniel W. E. Green